40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jan 4 12:46:27 2018, gautam, Update, ALS, Fiber ALS assay FiberALS.pdfFiberALS.svg.zip
    Reply  Thu Jan 11 14:22:40 2018, gautam, Update, ALS, Fiber ALS assay 
       Reply  Thu Jan 11 20:51:20 2018, gautam, Update, ALS, Fiber ALS assay IMG_5112.JPG
          Reply  Tue Jan 16 21:50:53 2018, gautam, Update, ALS, Fiber ALS assay 
             Reply  Thu Jan 18 00:35:00 2018, gautam, Update, ALS, Fiber ALS assay IR_ALS_20180118.pdfC2B4C1DD-6528-4067-9C13-6BD248629AD6.MOV
                Reply  Fri Jan 19 11:34:21 2018, gautam, Update, ALS, Fiber ALS assay IR_ALS_20180119.pdf
                   Reply  Fri Jan 19 23:04:11 2018, gautam, Update, ALS, Fiber ALS assay 
                      Reply  Wed Jan 24 00:33:31 2018, gautam, Update, ALS, Fiber ALS assay IMG_6866.JPG
                         Reply  Wed Jan 24 10:45:14 2018, gautam, Update, ALS, Fiber ALS assay 
                         Reply  Thu Jan 25 23:59:14 2018, gautam, Update, ALS, Fiber ALS assay 
                            Reply  Fri Jan 26 20:03:09 2018, gautam, Update, ALS, Fiber ALS assay 
                               Reply  Wed Jan 31 15:45:22 2018, gautam, Update, ALS, Fiber ALS assay IMG_6875.JPG
                                  Reply  Wed Jan 31 16:29:42 2018, gautam, Update, ALS, Modulation depths PMCmodDepth.pdfXPDH.pdf
                                  Reply  Wed Jan 31 16:33:53 2018, gautam, Update, ALS, ALS electronics at LSC rack 
                                     Reply  Wed Jan 31 22:32:11 2018, gautam, Update, ALS, ALS signal chain + power budget FiberALS_PowerBudget.pdfpreampProposed.pdf
                                        Reply  Thu Feb 1 01:24:56 2018, gautam, Update, ALS, D0902745 revamp underway PSLinterlock.png
                                           Reply  Fri Feb 2 00:26:34 2018, gautam, Update, ALS, D0902745 revamp underway 
                                              Reply  Fri Feb 2 13:16:55 2018, gautam, Update, ALS, ALS signals whitening switching ALS_whitening_switching.pdfALS_whitening_switching_works.pdf
                                                 Reply  Mon Feb 5 14:11:01 2018, gautam, Update, ALS, Huge harmonics in ALS channels ALSpowerlineHarmonics.pdf2A1B3AC4-4059-416A-AD88-8AB0431D7A55.jpeg
                                                    Reply  Mon Feb 5 22:57:28 2018, gautam, Update, ALS, Huge harmonics in ALS channels 
                                                       Reply  Tue Feb 6 11:13:26 2018, gautam, Update, ALS, Possible source of ground loop identified A68AF89C-E8A9-416D-BBD2-A1AD0A51E0B5.jpeg
                                                          Reply  Tue Feb 6 22:55:51 2018, gautam, Update, ALS, Possible source of ground loop identified 
                                                       Reply  Thu Feb 8 00:33:20 2018, gautam, Update, ALS, D990694 characterization / THD measurement plan D990694THD_trial.pdf
                                                          Reply  Thu Feb 8 01:27:16 2018, Koji, Update, ALS, D990694 characterization / THD measurement plan 
                                                          Reply  Fri Feb 9 13:37:44 2018, gautam, Update, ALS, THD measurement trial THD_trial.pdf
                                                       Reply  Thu Feb 8 12:00:09 2018, gautam, Update, ALS, D990694 is NOT differential receiving 
                                                          Reply  Thu Feb 8 13:13:14 2018, gautam, Update, ALS, D990694 pulled out IMG_5131.JPGIMG_5133.JPGD990694-B_cuttingPlan.pdf
                                                             Reply  Thu Feb 8 18:10:36 2018, gautam, Update, ALS, D990694 pulled out 
                                              Reply  Wed Feb 7 15:51:15 2018, gautam, Update, ALS, D0902745 revamp complete 6613CD37-5014-44AE-B1FE-6F6A8137DF62.jpeg
                                                 Reply  Fri Feb 16 01:34:40 2018, gautam, Update, ALS, D0902745 in-situ testing BeatMouthX_20180216.pdfIMG_5134.JPG
                                                    Reply  Tue Feb 20 23:08:27 2018, gautam, Update, ALS, D0902745 in-situ testing BeatMouthX_20180220_diffOut.pdf
                                                       Reply  Thu Feb 22 00:09:11 2018, gautam, Update, ALS, D0902745 in-situ testing daughterBoard_TF.pdfdaughterBoard_noise.pdfIMG_6906.JPG
                                                          Reply  Sun Feb 25 00:03:12 2018, gautam, Update, ALS, Daughter board prototyping schematic.pdfdaughterBoard_TF.pdfdaughterBoard_noise.pdfTEK00000.PNGdaughterBoard_noise.pdf
                                                             Reply  Mon Feb 26 20:55:56 2018, rana, Update, ALS, Daughter board prototyping 
                                                                Reply  Tue Feb 27 21:10:45 2018, gautam, Update, ALS, Daughter board testing IMG_6912.JPGIMG_6913.JPGIMG_6914.JPG
                                                                   Reply  Wed Feb 28 19:13:25 2018, gautam, Update, ALS, ADC test for differential receiving in c1lsc c1lsc_ADCtest.png
                                                                   Reply  Fri Mar 2 01:45:06 2018, gautam, Update, ALS, new look ALS electronics IMG_6916.JPG
                                                                      Reply  Mon Mar 5 17:27:34 2018, gautam, Update, ALS, new look ALS electronics - characterization BeatMouthX_20180305_diffOut.pdf
                                                                         Reply  Thu Mar 8 00:40:25 2018, gautam, Update, ALS, new look ALS electronics - characterization BeatMouthX_20180305_diffOut.pdf
                                                                            Reply  Thu Mar 8 19:38:37 2018, gautam, Update, ALS, digital unwhitening of daughter board TFfit.pdf
                                                                               Reply  Thu Mar 8 23:50:27 2018, gautam, Update, ALS, First look at new ALS electronics BeatMouth_OOL.pdf
                                                                                  Reply  Fri Mar 9 01:07:01 2018, gautam, Update, ALS, First look at new ALS electronics ALS_vs_POXnoise.pdf
                                                                                     Reply  Mon Mar 12 22:08:31 2018, gautam, Update, ALS, Noisy POX NoisyPOX_20180312.pdf
                                                                                        Reply  Mon Mar 12 23:57:31 2018, gautam, Update, ALS, Noisy POX NoisyPOXandPOY_20180312.pdfnoisyPOX_Jan2018.pdf
                                                                                           Reply  Mon Mar 19 15:02:29 2018, gautam, Update, ALS, Noisy MC sensing 
                                        Reply  Thu Feb 1 15:31:12 2018, gautam, Update, ALS, ALS signal chain + power budget 
                                        Reply  Wed Feb 7 10:16:26 2018, gautam, Update, ALS, ALS signal chain + power budget 
                      Reply  Thu Jan 25 13:18:41 2018, gautam, Update, ALS, Fiber ALS assay preampProposed.pdf
Message ID: 13552     Entry time: Tue Jan 16 21:50:53 2018     In reply to: 13534     Reply to this: 13557
Author: gautam 
Type: Update 
Category: ALS 
Subject: Fiber ALS assay 

With Johannes' help, I re-installed the box in the LSC electronics rack. In the end, I couldn't find a good solution to thermally insulate the inside of the box with foam - the 2U box is already pretty crowded with ~100m of cabling inside of it. So I just removed all the haphazardly placed foam and closed the box up for now. We can evaluate if thermal stability of the delay line is limiting us anywhere we care about and then think about what to do in this respect. This box is actually rather heavy with ~100m of cabling inside, and is right now mounted just by using the ears on the front - probably should try and implement a more robust mounting solution for the box with some rails for it to sit on.

I then restored all the cabling - but now, the delayed part of the split RF beat signal goes to the "RF in" input of the demod board, and the non-delayed part goes to the back-panel "LO" input. I also re-did the cabling at the PSL table, to connect the two ZHL3-A amplifier inputs to the IR beat PDs in the BeatMouth instead of the green BBPDs.

I didn't measure any power levels today, my plan was to try and get a quick ALS error signal spectrum - but looks like there is too much beat signal power available at the moment, the ADC inputs for both arm beat signals are overflowing often. The flat gain on the AS165 (=ALS X) and POP55 (=ALS Y) channels have been set to 0dB, but still the input signals seem way too large. The signals on the control room spectrum analyzer come from the "RF mon" ports on the demod board, and are marked as -23dBm. I looked at these peak heights with the end green beams locked to the arm cavities, as per the proposed new ALS scheme. Not sure how much cable loss we have from the LSC rack to the network analyzer, but assuming 3dB (which is the Google value for 100ft of RG58), and reading off the peak heights from the control room analyzer, I figure that we have ~0dBm of RF signal in the X arm. => I would expect ~3dBm of signal to the LO input. Both these numbers seem well within range of what the demod board is designed to handle so I'm not sure why we are saturating.

Note that the nominal (differential) I and Q demodulated outputs from the demod board come out of a backplane connector - but we seem to be using the front panel (single-ended) "MON" channels to acquire these signals. I also need to update my Fiber ALS diagram to indicate the power loss in cabling from the PSL table to the LSC electronics rack, expect it to be a couple of dB.

 

Quote:

After labeling cables I would disconnect, I pulled the box out of the LSC rack. Attachment #1 is a picture of the insides of the box - looks like it is indeed just two lengths of cabling. There was also some foam haphazardly stuck around inside - presumably an attempt at insulation/isolation.

Since I have the box out, I plan to measure the delay in each path, and also the signal attenuation. I'll also try and neaten the foam padding arrangement - Steve was showing me some foam we have, I'll use that. If anyone has comments on other changes that should be made / additional tests that should be done, please let me know.

20180111_2200: I'm running some TF measurements on the delay line box with the Agilent in the control room area (script running in tmux sesh on pianosa). Results will be uploaded later.

 

 

ELOG V3.1.3-